OpenStudy (anonymous):

What is the sum? 7/2 + (-3/5) 1. 4/3 2. 29/10 3. -21/10 4. 41/10

OpenStudy (anonymous):

\[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bd}\] \[\frac{7}{2}-\frac{3}{5}=\frac{7\times 5-3\times 2}{2\times 5}\] \[=\frac{35-6}{10}=...\]